4/3 Wednesday.
Jackup barge in place along bund wall in small craft anchorage with the Bhagwan Ocean, the Mildred and the Bhagwan Jackel coming and going with a dumb barge throughout the day between here and The Woomera, excavating the diamond Clinton main/Clinton Bypass. They have set an Anchor bouy in this area I have marked with a Way point 009.
Both mains required oil but only had 1ltr on board, used it and Ordered more oil by ESR.
MRF 5553 raised as requested by Willie- shore power connection on vessel loose?
5/3 Thursday
Brody tightened the shore power connection box on the vessel and greased inlet shaft. We connected shore power several times while berthed both port side to and STB side to, with no problems.
6/3 Friday
MRF 5560 raised as radar was not operational on departure from Marina and displayed messages like “No Radar”, “Unknown Radar Scanner”, & “Sources (invalid) 4G-991A. At 1600 the radar came back on line. Notified Brody he said we’ll just leave it for now see how it goes.
Some sort of anchor/mooring bouy located in our path from CP3-CP1 (track 3 the red one) coordinates S23 46.112 E151 10.145 created new way point 008.
7/3 Saturday
1100hrs No Radar again. Notified Brody and Willie, will have to wait till Monday to have it looked at. Radar came back online at 1740.
Life Jacket Demo preformed and then received feedback from Timmie Ramasray to the effect of “Master was not loud enough on the PA” (I am never softly spoken but I’m afraid I cannot compete with noise canceling headphones and ignorance) and “can we please do the demonstration prior to departure so it is less disruptive”
phrfff!!! Whatever they want!
8/3 Sunday
Gav on board to fix Radar. Preformed full system software update. Seems to be working ok. Please only select no to software update messages. Updates to be completed by Gav only.
Preformed drills. When anchoring the rope became jammed in the chain guide and we were unable to retrieve anchor. Willie had to board us in order to assist in retrieving the anchor manually to the point where the chain could be relocated into the chain guide. Takeaways NEVER let the anchor out past the 35mts of chain. Greame has requested the maintenance paint the chain in order for us to know when the rope is approaching.
9/3 Monday
stb main lost 70rpm. Brody on board to replace the waste gate actuator. All good.
3 monthly Audit by CoP marine personal Graeme on board. Outcome-the following issues where identified by the client. *Masters scribbling things out in the log book when they make a mistake rather than simply crossing it out neatly so the original entry can still be read and signing with initials beside it. *SDS folder contains documents which are over 5yrs old. *Defib pads have expired. Items raised by the crew *requested a wind sock for CP3.
CoP paramedic on board to change out defib pads.
received feedback in regards to life jacket demonstrations and sent an email to management highlighting CoP concerns as well as our concerns re pax boarding and disembarking whilst wearing headphones.
Received email notification that the anchor chain will be painted with indicators different colours in order to prevent us from accidentally splaying out past the chain. I suggested to Graeme that a small laminated card be made up to display beside the ‘Maday Distress call’ card on the helm.
10/3 Tuesday
Andy and I went through the SDS and identifyed all the out of date sheets and an email was sent through to Graeme listing all the replacements required.
Addendum No 12 in the SMS for you all to sign off on.
11/3 Wednesday
MRF’ed the decomposing superstructure as requested by Gene Mitchell. Also windscreen wiper blade.
Received instructions to up the cleaning regime re Corona Virus. Ref personal emails not sent to vessels.
12/3 Thursday
Andy went through our Safety Evacuation cards (in the backs of the seats) and I ordered 25 new ones.
13/3 Friday
New windscreen wiper arm fitted.
received new Safety Evacuation Cards and disposed of damaged ones.
16/3 Monday
The Curlew is commencing works on the CP3 and CP1 pontoons. They have our timetable and have been asked not to interrupt our schedule. They have been very good.
Stb side GPS monitor not displaying SOG, DEPTH, HDG, POS. Sent a picture to Brody and MRF 5592 was raised.
Canaipa has been assigned new scheduled runs, see new timetable.
17/3 Tuesday
Brody and Shaun on board to sort out the radar issue. Seems to be good now.
The Woomera has relocated in the Clinton swing basin.
Security has requested that we arrive at CP1 before 0430 for the 0500 run as a few people were there early this morning. Not sure if they are authorised to give us instructions but passing it on anyway.
18/3 Wednesday
Diva 8 dive Ops on barge poles at CP3 RoRo.
Received email notification to cease all life Jacket demonstrations re Corona virus till further notice.
GPS/Radar display started playing up again. No SOG, HDG, DEPTH. Also constantly getting NoGPS fix alarm. MRF raised 5598. Also rang Brody.
20/3 Friday
ordered engine oil by ESR.
New seating arrangement for CoP came into place today along with pax capping at 72. APLNG security is in charge of crowd control from the entry to the causeway and FLNE security from the turnstiles.
New hand sanitiser being supplied to be bunded into our existing pump bottles. I have requested by email 2 x sticky labels for the bottles along with the appropriate SDS sheet.
3 yearly maintenance checks submitted (nothing relevant).
21/3 Saturday
Excavation operation moving to and fro between the Clinton Diamond and the Clinton swing basin.
weekly, monthly and 3 monthly maintenance checks completed.
Port aft fire flap inoperable due to broken pin. MRF 5601 raised. Luke jimmied up the cable so it works now. ***awaiting to receive split pin in order to replace temporary fix (luke)***
23/3 Monday
On pre start checks the STB FM panel alarm would not reset. The panel on the dash reset but the alarm still sounded. Shaun replaced the panel. When you set the test off now you have to isolate and rest, it doesn’t do it automatically.
24/3 Tuesday
Received new 20ltrs engine oil and 5ltrs jet oil from maintenance this morning.
The Trojan has come on line across all three sites to transport all suspected COVID cases off the island. Had the first transfer from CP3 today. They will be tying up at Duffy’s berth and the patient will be escorted across to FLNE on the back deck.
A lot of ConocoPhillips staff now working from home though still bringing well over 100pax per day to the island? (Don’t know what happened to essentially personnel only?)
Pax numbers onboard are now reduced, monitor pax count and inform security
Refer to flagged emails.
Have been running with entry door windows (top ones) open as well as the back doors open in order to keep airflow through vessel. Would suggest continuing to do so
Liam, you might notice that the boat hook message never really got through....
Has anybody received the new fire flap stickers that were ordered 2 swings ago ?
Have a nice swing this is mine and Andy’s last one on BK. 😊